# SpokenWOZ: A Large-Scale Speech-Text Benchmark for Spoken Task-Oriented Dialogue Agents

### v5

@@ 1 Introduction
Dialogue system modeling, a classic research topic in the field of human-machine interaction, serves as an important application area (li2023pace, ; li2023vdialogue, ; si2023mining, ; zhao2023mmicl, ; si2022sclrai, ; si2023santa, ). Recently, task-oriented dialogue (TOD) systems (heck-etal-2020-trippy, ; yang2020ubar, ; cai2023unipcm, ) have attracted significant attention. These systems are designed to assist users in accomplishing specific goals, e.g., flight booking and restaurant reservation. Benefited from the previous well-designed data-collection pipeline, such as Wizard-of-Oz (DBLP:journals/tois/Kelley84, ), the TOD datasets (wen-etal-2017-network, ; budzianowski-etal-2018-multiwoz, ) can be constructed with low costs by writing from annotators, rather than being collected from realistic spoken conversations.
However, these TOD datasets constructed solely based on written texts may not accurately reflect the nuances of spoken conversations, leading to a gap between academic research and real-world spoken TOD scenarios. To advance the research on more realistic spoken TOD, various datasets have been introduced to simulate and model spoken conversations (hemphill-etal-1990-atis, ; henderson2014second, ; kim2021how, ). Despite these efforts, existing datasets still exhibit three notable weaknesses:
| Metric | DSTC2 | KVRET | M2M | MultiWOZ | ABCD | DSTC10 | SpokenWOZ |
@@
Data Scale: Previous spoken TOD datasets have been limited in terms of data scale, posing challenges for developing strong and more realistic systems. The pioneering ATIS dataset (hemphill-etal-1990-atis, ) consists of only 41 dialogues, which severely restricts the diversity and breadth of dialogues. While recent efforts, such as DSTC2 (henderson2014second, ) and DSTC10 (kim2021how, ), have renewed interest in modeling spoken TOD, they offer a modest improvement with 1,612 and 107 dialogues, respectively. In contrast, written TOD datasets offer a significantly larger number of dialogues, exemplified by MultiWOZ (budzianowski-etal-2018-multiwoz, ). Therefore, addressing this data scale limitation is crucial to advancing the spoken TOD systems.
@@ 2 Related Work
Written TOD datasets To push forward research in TOD modeling, various written TOD datasets have been proposed, ranging from human-to-machine (raux2005let, ) to more natural human-to-human dialogues (budzianowski-etal-2018-multiwoz, ), as well as from single-domain (wen-etal-2017-network, ) to more realistic multi-domain scenarios (DBLP:conf/aaai/RastogiZSGK20, ). Notable written TOD datasets include M2M (shah2018building, ), KVRET (eric2017key, ), MultiWOZ (budzianowski-etal-2018-multiwoz, ), SGD (rastogi2020towards, ), and ABCD (chen-etal-2021-action, ). Initially, these datasets focus on the single-domain dialogue. Subsequently, MultiWOZ and SGD are proposed to address more realistic multi-domain dialogue. Recently, numerous written TOD datasets attempt to simulate more realistic spoken conversations. For instance, RADDLE (peng-etal-2021-raddle, ) manually introduces various human-designed spoken noises to the written MultiWOZ. CGoDial (dai2022cgodial, ) adds spoken features to existing Chinese TOD datasets via crowd-sourcing, but the language and lack of audio data both limit the research on spoken TOD. Similarly, SSTOD (zhang-etal-2022-slot, ) introduces segment-by-segment interactions in simulated spoken language. Nonetheless, these studies concentrate solely on specific aspects of spoken TOD, thereby restricting a more holistic examination.
Spoken TOD datasets Several datasets have been proposed for modeling the spoken TOD. ATIS (hemphill-etal-1990-atis, ) focuses on single-domain TOD and only contains 41 dialogues. DSTC2 (henderson2014second, ) provides spoken corpora, but remains limited in scale. DSTC10 (kim2021how, ) revisits spoken TOD, providing 107 dialogues that include ASR hypotheses. EVI (spithourakis-etal-2022-evi, ) provides a transcript and ASR hypotheses, but only evaluates limited tasks, including enrolment (E), verification (V) and identification (I), ignoring more common tasks in TOD, such as querying and booking. To the best of our knowledge, SpokenWOZ represents the first large-scale speech-text dataset for TOD with annotation of the dialogue state and act.
@@ 3.1 Collection of Dialogue Audio
| acts |
inform / request / ack /select / recommend / nooffer / nobook / book edit / confirm / greet / bye / reqmore / wait / backchannel / thanks |
| slots |
address / postcode / phone / name / area / price range type / internet parking / stars / open hours / departure / destination leaveat / arriveby / ID / email / license plate number / bookpeople trainID / ticket price / booktime / department / bookday / day / bookstay |
@@ 3.2 Annotation of Dialogue
Quality Control. Before annotation, the annotators will return the audio in low quality. To ensure the quality of the annotations, we implement a three-step quality control process, including script checking based on rules, full inspection by annotators, and random inspection by us. We use scripts to identify simple errors such as missing annotations and return the dialogues for revision. Next, annotators conduct a full inspection and correct any wrong labels. Finally, we carry out a random inspection on 10% of the dialogues in SpokenWOZ. If the correct rate of the random inspection was less than 97% at turn-level, we would return this batch of dialogues for re-inspection by the annotators. Due to the strict quality control, SpokenWOZ can achieve a much higher quality than MultiWOZ, which contains over 40% of dialogue turns that need to be modified (eric-etal-2020-multiwoz, ).
@@ 4.2.2 Reasoning Slot Detection
To address the challenges presented by the casual nature of spoken language, TOD systems need to engage in reasoning to effectively complete slot updates. SpokenWOZ introduces three kinds of reasoning, including Temporal Reasoning (zhou-etal-2019-going, ), Mathematical Reasoning (patel-etal-2021-nlp, ), and Semantic Reasoning (sen-etal-2020-schema, ). More reasoning slot details can be found in Appendix A.1, including the name of the specific slot and how to construct the reasoning slots in dialogue.
@@ 5 Tasks & Settings
Dialogue State Tracking (DST). As highlighted in Section 4, the challenges in DST involve understanding noisy utterances, managing cross-turn slots, and handling reasoning slots. We use the joint goal accuracy (JGA) (budzianowski-etal-2018-multiwoz, ) as the metric, which measures the ratio of turns for which the value of each slot is correct.
@@
Following the previous TOD setting (budzianowski-etal-2018-multiwoz, ), we report the metrics of whether the system provides a correct entity (INFORM) and answers all the requested information (SUCCESS). A combined score (Comb) is computed by (INFORM+SUCCESS)×0.5+BLEU (papineni-etal-2002-bleu, ) as an overall quality measure.
@@ 6.1 Baselines
BERT+TripPy: TripPy (DBLP:conf/sigdial/HeckNLGLMG20, ) uses BERT (devlin-etal-2019-bert, ) as encoder and uses copy mechanisms to fill slots for DST, including span prediction, inform operations, and slot copy.
UBAR: UBAR (yang2021ubar, ) is acquired by fine-tuning the GPT-2 (Radford2019LanguageMA, ) on the sequence of the entire dialogue which is composed of user utterance, dialogue state, database result, system act, and system response.
GALAXY: GALAXY (DBLP:conf/aaai/HeDZWCLJYHSSL22, ) is a pre-trained conversation model based on UniLM (dong2019unified, ) that learns dialogue policy via semi-supervised learning. GALAXY keeps the same input and output format as UBAR to generate the dialogue state or system response.
SPACE: SPACE (DBLP:conf/sigir/HeDYSHSL22, ) is a pre-trained conversation model based on UniLM that takes into account both dialogue understanding and dialogue generation. SPACE keeps the same input and output format as UBAR to generate the dialogue state or system response.
@@
We utilize WavLM (DBLP:journals/jstsp/ChenWCWLCLKYXWZ22, ) as the speech encoder and implement the following dual-modal models:
SPACE+WavLM+TripPy: We use a bi-encoder framework and then concatenate the representation outputs from SPACE and WavLM, then use Transformer (DBLP:conf/nips/VaswaniSPUJGKP17, ) encoder as fusion layer to allow the interaction between the different modalities. We use the multi-modal outputs from the fusion layer as the representations for TripPy.
@@
SPACE+WavLM: As shown in Figure 3, we align the user-side word context and the corresponding audio pieces based on the annotations in SpokenWOZ. Then we add the user utterance embeddings from SPACE and the speech embeddings from WavLM as new user-side embeddings. During the inference, SPACE+WavLM uses dual-modal input to generate the state and response.
@@
| InstructGPT | 14.15 | 16.49 |
As LLM makes amazing progress on NLP tasks (DBLP:journals/corr/abs-2302-04023, ), we also evaluate LLM’s zero-shot performance. We use the prompts form Hudecek et al. (DBLP:journals/corr/abs-2304-06556, ) and Bang et al. (DBLP:journals/corr/abs-2302-04023, ) as reference.
ChatGPT: ChatGPT (GPT-3.5-turbo) is a conversational LLM (DBLP:conf/nips/Ouyang0JAWMZASR22, ), which has brought remarkable success on various zero-shot tasks.
InstructGPT: InstructGPT (text-davinci-003) (DBLP:conf/nips/Ouyang0JAWMZASR22, ) is a 175B LLM trained by instruction tuning and reinforcement learning.
@@ 6.2 Results and Discussion
Dual-modal TOD models is what you need. The significant size of the dialogues and audios in SpokenWOZ allows researchers to build large data-driven neural models given textual inputs and dual-modal inputs. To use dual-modal data in SpokenWOZ, we introduce several dual-modal baselines. Dual-modal methods achieve improved performance, showing the necessity of dual-modal methods in realistic scenarios. We observe that SPACE+WavLM outperforms SPACE+WavLM due to the speech-text alignment. We show Case Study in Appendix A.3 to confirm our claim.
@@
Dialogue state is the bottleneck of LLMs. As shown in 7 and 8, LLMs do not outperform supervised methods in DST and End-to-end Modeling. But, it is worth noting that LLMs achieve comparable performances in Policy Optimization task. It shows LLMs can complete the user request if provided the golden dialogue state and database results. However, when LLMs use their predicted dialogue state and database results in End-to-end Modeling, the performance is much lower. This suggests that LLM’s ability to accomplish dialogue goals depends on the predictions of the dialogue state. To explore the reasons for the poor performance of LLMs, we provide an analysis in Appendix A.9. We find the main reason in DST is that the hallucination phenomenon (marie-fujita-2020-synthesizing, ) is very serious, i.e., LLMs generate erroneous results at slots that are not involved in the conversation. Meanwhile, LLMs are sensitive to noisy utterances, e.g., LLMs tend to directly copy the noisy word in the utterance as result, which may be due to the inability of LLMs to perceive the speech information.
@@ A.1.1 Conversation Details
| Temporal Reasoning | leaveat arriveby booktime / day bookday |
| Mathematical Reasoning | bookpeople bookstay |
| Semantic Reasoning | type area internet department parking |
@@ A.2.1 DST Baselines
InstructGPT:
InstructGPT (text-davinci-003) [24] is a 175B LLM trained by reinforcement learning with human feedback and instruction tuning.
@@ A.2.2 Response Generation Baselines
SPACE+WavLM
SPACE+WavLM adds the text embeddings from SPACE and the corresponding embeddings from WavLM as new user-side embeddings. During the inference, the model uses dual-modal input to generate dialogue state, act, and final system response.
@@ A.9.1 Prompts for LLMs
For response generation tasks, we follow the prompt as Hudecek et al. [18]. We use the dialogue state to query the database provided to get the entity. For the Policy Optimization task, we use the ground truth dialogue state to query the provided database. For End-to-end Modeling, we use the predicted dialogue state generated from ChatGPT or InstructGPT to query the provided database. We use the contents of the database and the generated system response to match the strings and complete the de-lexicalization process. The de-lexicalized response will be used to calculate INFORM and SUCCESS. We list a prompt example for a response generation task in Table 21.
@@ A.9.2 Analysis on DST
As shown in the section Experiments, the performances of LLMs do not surpass that of supervised small models and show a noticeable gap compared to the supervised generative dual-modal baselines in the DST task. However, it is worth noting that LLMs slightly outperform the BERT+TripPy when cross-turn slots are not taken into account, showing the potential capabilities of LLMs. Meanwhile, we find that the main reason for the poor results of LLMs is that the hallucination phenomenon [23] is very serious, e.g., LLMs often generate additional results that do not fit the dialogue utterances. When we measure only the slots where the ground truth value is not “none”, we find that ChatGPT achieves JGA of 30.81 and InstructGPT achieves JGA of 34.42. Compared to the standard JGA, ChatGPT and InstructGPT improve their scores by 17.06 and 20.27 respectively. This indicates that LLMs generate a large number of erroneous results at slots that are not involved in the conversation, i.e., the hallucination phenomenon is very serious. We show a case in Table 24.
@@
| (InstructGPT: Hotel-Name = lavelle lodge) |
@@ E.5 Uses
We provide links to the papers of all the baseline models on the leaderboard444https://spokenwoz.github.io/SpokenWOZ-github.io/ we built.
@@ E.6 Distribution
SpokenWOZ dataset and codebases for reproducing the experiments are available at: https://spokenwoz.github.io/SpokenWOZ-github.io/.
@@
The dataset is now available at: https://spokenwoz.github.io/SpokenWOZ-github.io/.

### v6

@@ 1 Introduction
Dialogue system modeling, a classic research topic in the field of human-machine interaction, serves as an important application area [22, 21, 35, 44, 36, 34]. Recently, task-oriented dialogue (TOD) systems [14, 41, 3] have attracted significant attention. These systems are designed to assist users in accomplishing specific goals, e.g., flight booking and restaurant reservation. Benefited from the previous well-designed data-collection pipeline, such as Wizard-of-Oz [19], the TOD datasets [39, 2] can be constructed with low costs by writing from annotators, rather than being collected from realistic spoken conversations.
However, these TOD datasets constructed solely based on written texts may not accurately reflect the nuances of spoken conversations, leading to a gap between academic research and real-world spoken TOD scenarios. To advance the research on more realistic spoken TOD, various datasets have been introduced to simulate and model spoken conversations [16, 17, 20]. Despite these efforts, existing datasets still exhibit three notable weaknesses:
| Metric | DSTC2 | KVRET | M2M | MultiWOZ | ABCD | DSTC10 | SpokenWOZ∗ |
@@
Data Scale: Previous spoken TOD datasets have been limited in terms of data scale, posing challenges for developing strong and more realistic systems. The pioneering ATIS dataset [16] consists of only 41 dialogues, which severely restricts the diversity and breadth of dialogues. While recent efforts, such as DSTC2 [17] and DSTC10 [20], have renewed interest in modeling spoken TOD, they offer a modest improvement with 1,612 and 107 dialogues, respectively. In contrast, written TOD datasets offer a significantly larger number of dialogues, exemplified by MultiWOZ [2]. Therefore, addressing this data scale limitation is crucial to advancing the spoken TOD systems.
@@ 2 Related Work
Written TOD datasets To push forward research in TOD modeling, various written TOD datasets have been proposed, ranging from human-to-machine [31] to more natural human-to-human dialogues [2], as well as from single-domain [39] to more realistic multi-domain scenarios [29]. Notable written TOD datasets include M2M [33], KVRET [11], MultiWOZ [2], SGD [30], and ABCD [4]. Initially, these datasets focus on the single-domain dialogue. Subsequently, MultiWOZ and SGD are proposed to address more realistic multi-domain dialogue. Recently, numerous written TOD datasets attempt to simulate more realistic spoken conversations. For instance, RADDLE [27] manually introduces various human-designed spoken noises to the written MultiWOZ. CGoDial [6] adds spoken features to existing Chinese TOD datasets via crowd-sourcing, but the language and lack of audio data both limit the research on spoken TOD. Similarly, SSTOD [43] introduces segment-by-segment interactions in simulated spoken language. Nonetheless, these studies concentrate solely on specific aspects of spoken TOD, thereby restricting a more holistic examination.
Spoken TOD datasets Several datasets have been proposed for modeling the spoken TOD. ATIS [16] focuses on single-domain TOD and only contains 41 dialogues. DSTC2 [17] provides spoken corpora, but remains limited in scale. DSTC10 [20] revisits spoken TOD, providing 107 dialogues that include ASR hypotheses. EVI [37] provides a transcript and ASR hypotheses, but only evaluates limited tasks, including enrolment (E), verification (V) and identification (I), ignoring more common tasks in TOD, such as querying and booking. To the best of our knowledge, SpokenWOZ represents the first large-scale speech-text dataset for TOD with annotation of the dialogue state and act.
@@ 3.1 Collection of Dialogue Audio
| acts | inform∗ / request∗ / ack∗ /select123 / recommend123 / nooffer123 / nobook125 / book125 edit8 / confirm8 / greet / bye / reqmore / wait / backchannel / thanks |
| slots | address12367 / postcode 12367 / phone 123678 / name12348 / area123 / price range123 type123 / internet2 parking2 / stars2 / open hours3 / departure45 / destination45 leaveat45 / arriveby45 / ID8 / email8 / license plate number8 / bookpeople125 trainID5 / ticket price5 / booktime1 / department7 / bookday12 / day5 / bookstay12 |
@@ 3.2 Annotation of Dialogue
Quality Control. Before annotation, the annotators will return the audio in low quality. To ensure the quality of the annotations, we implement a three-step quality control process, including script checking based on rules, full inspection by annotators, and random inspection by us. We use scripts to identify simple errors such as missing annotations and return the dialogues for revision. Next, annotators conduct a full inspection and correct any wrong labels. Finally, we carry out a random inspection on 10% of the dialogues in SpokenWOZ. If the correct rate of the random inspection was less than 97% at turn-level, we would return this batch of dialogues for re-inspection by the annotators. Due to the strict quality control, SpokenWOZ can achieve a much higher quality than MultiWOZ, which contains over 40% of dialogue turns that need to be modified [10].
@@ 4.2.2 Reasoning Slot Detection
To address the challenges presented by the casual nature of spoken language, TOD systems need to engage in reasoning to effectively complete slot updates. SpokenWOZ introduces three kinds of reasoning, including Temporal Reasoning [45], Mathematical Reasoning [26], and Semantic Reasoning [32]. More reasoning slot details can be found in Appendix A.1, including the name of the specific slot and how to construct the reasoning slots in dialogue.
@@ 5 Tasks & Settings
Dialogue State Tracking (DST). As highlighted in Section 4, the challenges in DST involve understanding noisy utterances, managing cross-turn slots, and handling reasoning slots. We use the joint goal accuracy (JGA) [2] as the metric, which measures the ratio of turns for which the value of each slot is correct.
@@ 6.1 Baselines
BERT+TripPy: TripPy [15] uses BERT [7] as encoder and uses copy mechanisms to fill slots for DST, including span prediction, inform operations, and slot copy.
UBAR: UBAR [40] is acquired by fine-tuning the GPT-2 [28] on the sequence of the entire dialogue which is composed of user utterance, dialogue state, database result, system act, and system response.
GALAXY: GALAXY [13] is a pre-trained conversation model based on UniLM [9] that learns dialogue policy via semi-supervised learning. GALAXY keeps the same input and output format as UBAR to generate the dialogue state or system response.
SPACE: SPACE [12] is a pre-trained conversation model based on UniLM that takes into account both dialogue understanding and dialogue generation. SPACE keeps the same input and output format as UBAR to generate the dialogue state or system response.
@@
We utilize WavLM [5] as the speech encoder and implement the following dual-modal models:
SPACE+WavLM+TripPy: We use a bi-encoder framework and then concatenate the representation outputs from SPACE and WavLM, then use Transformer [38] encoder as fusion layer to allow the interaction between the different modalities. We use the multi-modal outputs from the fusion layer as the representations for TripPy.
@@
SPACE+WavLMalign: As shown in Figure 3, we align the user-side word context and the corresponding audio pieces based on the annotations in SpokenWOZ. Then we add the user utterance embeddings from SPACE and the speech embeddings from WavLM as new user-side embeddings. During the inference, SPACE+WavLMalign uses dual-modal input to generate the state and response.
@@
| InstructGPT003 | 14.15 | 16.49 |
As LLM makes amazing progress on NLP tasks [1], we also evaluate LLM’s zero-shot performance. We use the prompts form Hudecek et al. [18] and Bang et al. [1] as reference.
ChatGPT: ChatGPT (GPT-3.5-turbo) is a conversational LLM [24], which has brought remarkable success on various zero-shot tasks.
InstructGPT003: InstructGPT003 (text-davinci-003) [24] is a 175B LLM trained by instruction tuning and reinforcement learning.
@@ 6.2 Results and Discussion
Dual-modal TOD models is what you need. The significant size of the dialogues and audios in SpokenWOZ allows researchers to build large data-driven neural models given textual inputs and dual-modal inputs. To use dual-modal data in SpokenWOZ, we introduce several dual-modal baselines. Dual-modal methods achieve improved performance, showing the necessity of dual-modal methods in realistic scenarios. We observe that SPACE+WavLMalign outperforms SPACE+WavLM due to the speech-text alignment. We show Case Study in Appendix A.3 to confirm our claim.
@@
Dialogue state is the bottleneck of LLMs. As shown in 7 and 8, LLMs do not outperform supervised methods in DST and End-to-end Modeling. But, it is worth noting that LLMs achieve comparable performances in Policy Optimization task. It shows LLMs can complete the user request if provided the golden dialogue state and database results. However, when LLMs use their predicted dialogue state and database results in End-to-end Modeling, the performance is much lower. This suggests that LLM’s ability to accomplish dialogue goals depends on the predictions of the dialogue state. To explore the reasons for the poor performance of LLMs, we provide an analysis in Appendix A.9. We find the main reason in DST is that the hallucination phenomenon [23] is very serious, i.e., LLMs generate erroneous results at slots that are not involved in the conversation. Meanwhile, LLMs are sensitive to noisy utterances, e.g., LLMs tend to directly copy the noisy word in the utterance as result, which may be due to the inability of LLMs to perceive the speech information.
@@ A.1.1 Conversation Details
| Temporal Reasoning | leaveat4,5 arriveby4,5 booktime1 / day5 bookday1,2 |
| Mathematical Reasoning | bookpeople1,2,5 bookstay2 |
| Semantic Reasoning | type1,2,3 area1,2,3 internet2 department6 parking2 |
@@ A.2.1 DST Baselines
InstructGPT003:
InstructGPT003 (text-davinci-003) [24] is a 175B LLM trained by reinforcement learning with human feedback and instruction tuning.
@@ A.2.2 Response Generation Baselines
SPACE+WavLMaligned
SPACE+WavLMaligned adds the text embeddings from SPACE and the corresponding embeddings from WavLM as new user-side embeddings. During the inference, the model uses dual-modal input to generate dialogue state, act, and final system response.
@@ A.9.1 Prompts for LLMs
For response generation tasks, we follow the prompt as Hudecek et al. [18]. We use the dialogue state to query the database provided to get the entity. For the Policy Optimization task, we use the ground truth dialogue state to query the provided database. For End-to-end Modeling, we use the predicted dialogue state generated from ChatGPT or InstructGPT003 to query the provided database. We use the contents of the database and the generated system response to match the strings and complete the de-lexicalization process. The de-lexicalized response will be used to calculate INFORM and SUCCESS. We list a prompt example for a response generation task in Table 21.
@@ A.9.2 Analysis on DST
As shown in the section Experiments, the performances of LLMs do not surpass that of supervised small models and show a noticeable gap compared to the supervised generative dual-modal baselines in the DST task. However, it is worth noting that LLMs slightly outperform the BERT+TripPy when cross-turn slots are not taken into account, showing the potential capabilities of LLMs. Meanwhile, we find that the main reason for the poor results of LLMs is that the hallucination phenomenon [23] is very serious, e.g., LLMs often generate additional results that do not fit the dialogue utterances. When we measure only the slots where the ground truth value is not “none”, we find that ChatGPT achieves JGA of 30.81 and InstructGPT003 achieves JGA of 34.42. Compared to the standard JGA, ChatGPT and InstructGPT003 improve their scores by 17.06 and 20.27 respectively. This indicates that LLMs generate a large number of erroneous results at slots that are not involved in the conversation, i.e., the hallucination phenomenon is very serious. We show a case in Table 24.
@@
| (InstructGPT003: Hotel-Name = lavelle lodge) |
@@ E.5 Uses
We provide links to the papers of all the baseline models on the leaderboard444https://spokenwoz.github.io/ we built.
@@ E.6 Distribution
SpokenWOZ dataset and codebases for reproducing the experiments are available at: https://spokenwoz.github.io/.
@@
The dataset is now available at: https://spokenwoz.github.io/.